Find (g o f)(3) if g(x) = 3x and f(x) = x - 3 Unfortunately, I can't solve it.
Answer by lwsshak3(11628)

Find (g o f)(3) if g(x) = 3x and f(x) = x - 3
**
(g o f)=g[f(x)]=3(x-3)=3x-9
(g o f)(3)=9-9=0